Happy Ten produces sports socks. The company has fixed expenses of $85,000 and variable expenses of $0.85 per package. Each package sells for $1.70.

Begin by identifying the formula to compute the contribution margin per package. Then compute the contribution margin per package.

1.

Compute the contribution margin per package and the contribution margin ratio.

2.

Find the breakeven point in units and in dollars.

3.

Find the number of packages Happy Ten needs to sell to earn 25,500
operating income.
Answer #1

Selling price per unit - Variable cost per unit = Contribution margin per unit

= 1.70 - 0.85

= 0.85

Contribution margin per unit / Selling price per unit = Contribution margin ratio

= 0.85/1.70

= 50%

Fixed cost / Contribution margin per unit = Break even point in units

85,000/0.85

= 100,000

Fixed cost / Contribution margin ratio = Breakeven point in dollars

= 85,000/50%

= 170,000

(Fixed cost + Target profit)/Contribution margin per unit = Sales needed

= (85,000+25,500)/0.85

= 130,000